# CausalPre: Scalable and Effective Data Pre-processing for Causal Fairness

### v1

@@ IV-B Refining the Fair DAG
we observe that only the conditional distribution of $Y$ differs, changing from $\mathbb{P}[Y\mid\Pi_{Y}]$ to $\mathbb{P}[Y\mid\Pi_{Y}\setminus(\mathcal{S}\cup\mathcal{I})]$, while the joint distribution over attributes in $\mathcal{V}\setminus\{Y\}$ remains unchanged. This insight allows us to abstract away the internal dependencies among non-label attributes and treat them collectively as a black box. Accordingly, when using the DAG as a processing guideline, it suffices to focus on the simplified structure in Figure 2(b), where the gray block represents the black-boxed relationships, rather than the full structure in Figure 2(a). Within this reduced view, we only need to identify the parent set of the label attribute (highlighted in the dashed orange box) and eliminate directed edges from sensitive or inadmissible attributes to the label for fairness issues.
@@
Building on this insight, we further refine the computation of $\mathbb{P}_{\mathcal{G}^{\prime}}$ by greedily selecting all attributes in $\mathcal{V}\setminus(\mathcal{S}\cup\mathcal{I}\cup{Y})$ as parents of $Y$. This refinement serves as an approximation that reduces the required structural guidance from Figure 2(b) to the simpler Figure 2(c), where the dashed orange box is no longer needed. Although this approximation may introduce additional edges directed to $Y$, the resulting factors are admissible by definition and thus acceptable for decision making. Empirically, we observe that the number of such additional edges is typically small, since most attributes in real-world datasets already act as decision-relevant factors.
@@ IV-D Accelerating Clique Selection
1 Initialize $\mathscr{C}_{\text{init}}\leftarrow\varnothing$, $\mathcal{U}\leftarrow\mathcal{X}$, $\mathcal{E}^{\prime}\leftarrow\mathcal{E}$, FLAG $\leftarrow$ TRUE;
2 Set number of clusters $r\leftarrow\left\lceil\frac{d-m-1}{k}\right\rceil$;
3
4while $|\mathscr{C}_{\text{init}}|<r$ do
5  Select edge $\langle i,j\rangle$ with the smallest $\mathcal{E}_{i,j}^{\prime}$, where $\mathcal{E}_{i,j}^{\prime}\in\mathcal{E}^{\prime}$;
6  for $e\in\{i,j\}$ do
7     if $X_{e}\in\mathcal{U}$ and $|\mathscr{C}_{\text{init}}|<r$ then
8        $\mathscr{C}_{\text{init}}\leftarrow\mathscr{C}_{\text{init}}\cup\{\{X_{e}\}\}$; $\mathcal{U}\leftarrow\mathcal{U}\setminus\{X_{e}\}$;
9
10
11 Remove $\mathcal{E}^{\prime}_{i,j}$ from $\mathcal{E}^{\prime}$;
12
13
14while $\mathcal{U}\neq\emptyset$ do
15  Identify $\langle X_{u},\mathcal{C}_{i}\rangle$ pair with the largest $\Delta$ value, where $\Delta$ is calculated by Equation (13), $X_{u}\in\mathcal{U}$, $\mathcal{C}_{i}\in\mathscr{C}_{\text{init}}$, and $|\mathcal{C}_{i}|<k{+}m$ if FLAG else $k$;
16  Update FLAG $\leftarrow$ FALSE if $|\mathcal{C}_{i}|>k$;
17  Update $\mathscr{C}_{\text{init}}$ by doing $\mathcal{C}_{i}\leftarrow\mathcal{C}_{i}\cup\{X_{u}\}$;
18  Remove $X_{i}$ from $\mathcal{U}$;
19
20 return $\mathscr{C}_{\text{init}}$;
@@ 18  Remove $X_{i}$ from $\mathcal{U}$;
1
[…]
3 Initialize $\mathscr{F}\leftarrow\varnothing$, $\mathscr{C}\leftarrow\{\{\mathcal{C}_{\text{active}}\}\}$, $\mathscr{U}\leftarrow\mathscr{C}_{\text{init}}\setminus\{\mathcal{C}_{\text{active}}\}$;
4
5while $\mathscr{U}\neq\varnothing$ do
6  Identify $\langle\mathcal{C}_{\text{active}},\mathcal{C}_{\text{opt}}\rangle$ with the largest $\Delta_{m}$, where $\mathcal{C}_{\text{active}}\in\mathscr{U}$, $\mathcal{C}_{\text{opt}}\in\mathscr{C}$, and $\Delta_{m}$ is from Equation (14);
7  $\mathcal{F}_{\text{active},\text{opt}}\leftarrow$ top-$m$ attributes from $\mathcal{C}_{\text{opt}}$ for $\Delta_{m}$;
8  Update $\mathscr{C}\leftarrow\mathscr{C}\cup\{\mathcal{C}_{\text{active}}\cup\mathcal{F}_{\text{active},\text{opt}}\}$;
9  Remove $\mathcal{C}_{\text{active}}$ from $\mathscr{U}$;
10
11
12return $\mathscr{C}$;
@@ IV-E The Complete Framework
2
3Initialize $\mathcal{E}\leftarrow\varnothing$;
4 for each pair of attributes $\langle X_{i},X_{j}\rangle\in\mathcal{V}$ do
5  Compute MI $\mathcal{E}_{i,j}=I[X_{i};X_{j}]$ and add it to $\mathcal{E}$;
6
[…]
8 $\mathscr{C}\leftarrow\textsc{CliqueExtension}(\mathcal{V},\mathscr{C}_{\text{init}},m)$;
9 $\mathcal{F}_{Y}\leftarrow$ Select (k+m1)(k{+}m{-}1) attributes from $\mathcal{V}\setminus(\mathcal{S}\cup\mathcal{I}\cup\{Y\})$ with strongest relationship to $Y$;
10 Update $\mathscr{C}\leftarrow\mathscr{C}\cup\{\mathcal{F}_{Y}\cup\{Y\}\}$;
11
@@ 10 Update $\mathscr{C}\leftarrow\mathscr{C}\cup\{\mathcal{F}_{Y}\cup\{Y\}\}$;
12
13Set $n\leftarrow|\mathcal{D}|$ ;
14 for $i\leftarrow 1$ to $r+1$ do
15  if $i=1$ then
16     Initialize $\mathcal{D}^{\prime}$ by sampling $n$ values for attributes in $\mathcal{C}_{1}$ from $\mathbb{P}[\mathcal{C}_{1}]$;
17
18 else
19     $\mathcal{C}_{i-1},\mathcal{C}_{i}\leftarrow$ the (i1i{-}1)-th, $i$-th clique in $\mathscr{C}$;
20     Identify separator $\mathcal{F}_{i-1,i}\leftarrow\mathcal{C}_{i-1}\cap\mathcal{C}_{i}$;
21     Sample $n$ values for $\mathcal{C}_{i}\setminus\mathcal{F}_{i-1,i}$ from $\mathbb{P}[\mathcal{C}_{i}\mid\mathcal{F}_{i-1,i}]$, and fill the corresponding columns in $\mathcal{D}^{\prime}$;
22
23
24
25return $\mathcal{D}^{\prime}$;
@@ V-A Setup
We implement the proposed CausalPre framework in Python. All experiments are conducted on a machine with two Xeon(R) Gold 6326@2.90 GHz CPUs and 256GB of DRAM.
@@
Figure 4: End-to-end performance on real-world datasets. The grey area indicates invalid regions.
@@ V-F Scalability Analysis
To assess scalability, we benchmark CausalPre and baseline methods on synthetic datasets containing 60 million records and 10-70 attributes, where each attribute has a domain size of 3-14. Figure 10 presents the memory usage and runtime of CausalPre and Cap-MF under varying attribute dimensionalities. Other baselines could not complete even the smallest dataset and are therefore omitted from this analysis. As shown, both memory usage and runtime for CausalPre increase steadily with the number of attributes. In contrast, Cap-MF exhibits exponential growth in both metrics and exhausts memory limits at 20 attributes. To further assess the end-to-end performance of CausalPre on high-dimensional data, we evaluate its utility and ROD on these synthetic datasets, with results shown in Figure 10. Since the datasets with different numbers of attributes are independently generated, we report relative utility and relative ROD, which are computed by normalizing against the corresponding values from “Original”. Cap-MF is excluded from this analysis as it does not scale effectively and cannot be applied to most of the evaluated cases. The results indicate that CausalPre preserves more than 99% of the original utility while reducing discriminatory effects by over 50% across all datasets through fairness-aware pre-processing.

### v2

@@ IV-B Refining the Fair DAG
we observe that only the conditional distribution of $Y$ differs, changing from $\mathbb{P}[Y\mid\Pi_{Y}]$ to $\mathbb{P}[Y\mid\Pi_{Y}\setminus(\mathcal{S}\cup\mathcal{I})]$, while the joint distribution over attributes in $\mathcal{V}\setminus\{Y\}$ remains unchanged. This insight allows us to abstract away the internal dependencies among non-label attributes and treat them collectively as a black box. Accordingly, when using the DAG as a processing guideline, it suffices to focus on the simplified structure in Figure 2b, where the gray block represents the black-boxed relationships, rather than the full structure in Figure 2a. Within this reduced view, we only need to identify the parent set of the label attribute (highlighted in the dashed orange box) and eliminate directed edges from sensitive or inadmissible attributes to the label for fairness issues.
@@
Building on this insight, we further refine the computation of $\mathbb{P}_{\mathcal{G}^{\prime}}$ by greedily selecting all attributes in $\mathcal{V}\setminus(\mathcal{S}\cup\mathcal{I}\cup{Y})$ as parents of $Y$. This refinement serves as an approximation that reduces the required structural guidance from Figure 2b to the simpler Figure 2c, where the dashed orange box is no longer needed. Although this approximation may introduce additional edges directed to $Y$, the resulting factors are admissible by definition and thus acceptable for decision making. Empirically, we observe that the number of such additional edges is typically small, since most attributes in real-world datasets already act as decision-relevant factors.
@@ IV-D Accelerating Clique Selection
Initialize $\mathscr{C}_{\text{init}}\leftarrow\varnothing$, $\mathcal{U}\leftarrow\mathcal{X}$, $\mathcal{E}^{\prime}\leftarrow\mathcal{E}$, FLAG $\leftarrow$ TRUE;
Set number of clusters $r\leftarrow\left\lceil\frac{d-m-1}{k}\right\rceil$;
while $|\mathscr{C}_{\text{init}}|<r$ do
  Select edge $\langle i,j\rangle$ with the smallest $\mathcal{E}_{i,j}^{\prime}$, where $\mathcal{E}_{i,j}^{\prime}\in\mathcal{E}^{\prime}$;
  for $e\in\{i,j\}$ do
     if $X_{e}\in\mathcal{U}$ and $|\mathscr{C}_{\text{init}}|<r$ then
        $\mathscr{C}_{\text{init}}\leftarrow\mathscr{C}_{\text{init}}\cup\{\{X_{e}\}\}$; $\mathcal{U}\leftarrow\mathcal{U}\setminus\{X_{e}\}$;
     end if
  end for
 Remove $\mathcal{E}^{\prime}_{i,j}$ from $\mathcal{E}^{\prime}$;
end while
while $\mathcal{U}\neq\emptyset$ do
  Identify $\langle X_{u},\mathcal{C}_{i}\rangle$ pair with the largest $\Delta$ value, where $\Delta$ is calculated by Equation (13), $X_{u}\in\mathcal{U}$, $\mathcal{C}_{i}\in\mathscr{C}_{\text{init}}$, and $|\mathcal{C}_{i}|<k{+}m$ if FLAG else $k$;
  Update FLAG $\leftarrow$ FALSE if $|\mathcal{C}_{i}|>k$;
  Update $\mathscr{C}_{\text{init}}$ by doing $\mathcal{C}_{i}\leftarrow\mathcal{C}_{i}\cup\{X_{u}\}$;
  Remove $X_{i}$ from $\mathcal{U}$;
end while
return $\mathscr{C}_{\text{init}}$;
@@
$\mathcal{C}_{\text{active}}\leftarrow$ The clique in $\mathscr{C}_{\text{init}}$ with size (k+m)(k{+}m);
Initialize $\mathscr{F}\leftarrow\varnothing$, $\mathscr{C}\leftarrow\{\{\mathcal{C}_{\text{active}}\}\}$, $\mathscr{U}\leftarrow\mathscr{C}_{\text{init}}\setminus\{\mathcal{C}_{\text{active}}\}$;
while $\mathscr{U}\neq\varnothing$ do
  Identify $\langle\mathcal{C}_{\text{active}},\mathcal{C}_{\text{opt}}\rangle$ with the largest $\Delta_{m}$, where $\mathcal{C}_{\text{active}}\in\mathscr{U}$, $\mathcal{C}_{\text{opt}}\in\mathscr{C}$, and $\Delta_{m}$ is from Equation (14);
  $\mathcal{F}_{\text{active},\text{opt}}\leftarrow$ top-$m$ attributes from $\mathcal{C}_{\text{opt}}$ for $\Delta_{m}$;
  Update $\mathscr{C}\leftarrow\mathscr{C}\cup\{\mathcal{C}_{\text{active}}\cup\mathcal{F}_{\text{active},\text{opt}}\}$;
  Remove $\mathcal{C}_{\text{active}}$ from $\mathscr{U}$;
end while
return $\mathscr{C}$;
@@ IV-E The Complete Framework
Initialize $\mathcal{E}\leftarrow\varnothing$;
for each pair of attributes $\langle X_{i},X_{j}\rangle\in\mathcal{V}$ do
  Compute MI $\mathcal{E}_{i,j}=I[X_{i};X_{j}]$ and add it to $\mathcal{E}$;
end for
$\mathscr{C}_{\text{init}}\leftarrow\textsc{CliqueInitialization}(\mathcal{V},\mathcal{E},k,m)$;
$\mathscr{C}\leftarrow\textsc{CliqueExtension}(\mathcal{V},\mathscr{C}_{\text{init}},m)$;
$\mathcal{F}_{Y}\leftarrow$ Select (k+m1)(k{+}m{-}1) attributes from $\mathcal{V}\setminus(\mathcal{S}\cup\mathcal{I}\cup\{Y\})$ with strongest relationship to $Y$;
Update $\mathscr{C}\leftarrow\mathscr{C}\cup\{\mathcal{F}_{Y}\cup\{Y\}\}$;
@@
Set $n\leftarrow|\mathcal{D}|$ ;
for $i\leftarrow 1$ to $r+1$ do
  if $i=1$ then
     Initialize $\mathcal{D}^{\prime}$ by sampling $n$ values for attributes in $\mathcal{C}_{1}$ from $\mathbb{P}[\mathcal{C}_{1}]$;
 else
     $\mathcal{C}_{i-1},\mathcal{C}_{i}\leftarrow$ the (i1i{-}1)-th, $i$-th clique in $\mathscr{C}$;
     Identify separator $\mathcal{F}_{i-1,i}\leftarrow\mathcal{C}_{i-1}\cap\mathcal{C}_{i}$;
     Sample $n$ values for $\mathcal{C}_{i}\setminus\mathcal{F}_{i-1,i}$ from $\mathbb{P}[\mathcal{C}_{i}\mid\mathcal{F}_{i-1,i}]$, and fill the corresponding columns in $\mathcal{D}^{\prime}$;
  end if
end for
return $\mathcal{D}^{\prime}$;
@@ V-A Setup
We implement the proposed CausalPre framework in Python [^2] Available at https://github.com/iamzhengying/CausalPre.git. All experiments are conducted on a machine with two Xeon(R) Gold 6326@2.90 GHz CPUs and 256GB of DRAM.
@@
Figure 4: End-to-end performance on real-world datasets. The gray shaded area indicates invalid regions.
@@ V-F Scalability Analysis
To assess scalability, we benchmark CausalPre and baseline methods on synthetic datasets containing 60 million records and 10–70 attributes, where each attribute has a domain size of 3–14. Figure 10 presents the memory usage and runtime of CausalPre and Cap-MF under varying attribute dimensionalities. Other baselines could not complete even the smallest dataset and are therefore omitted from this analysis. As shown, both memory usage and runtime for CausalPre increase steadily with the number of attributes. In contrast, Cap-MF exhibits exponential growth in both metrics and exhausts memory limits at 20 attributes. To further assess the end-to-end performance of CausalPre on high-dimensional data, we evaluate its utility and ROD on these synthetic datasets, with results shown in Figure 10. Since the datasets with different numbers of attributes are independently generated, we report relative utility and relative ROD, which are computed by normalizing against the corresponding values from “Original”. Cap-MF is excluded from this analysis as it does not scale effectively and cannot be applied to most of the evaluated cases. The results indicate that CausalPre preserves more than 99% of the original utility while reducing discriminatory effects by over 50% across all datasets through fairness-aware pre-processing.